At Twilight Coffee Shop, Manager Bella buys only half and half (12%)milk and skim (0%fat) milk to simplify inventory.
She is rushing out of town and need to quickly leave instructions for her employees on how to obtain 2% fat milk for customers.
A) For one set of instructions, she would like to end with 1 gallon of 2% fat milk. How much half and half and skim milk must she mix together to obtain 1 gallon of 2% fat milk?

Answer by jrfrunner(365)

let H=amount of Half and half
Let S = amount of skim milk
--
Given: information for two different equations (total amount, fat content)
1) total: H+S=1gallon
2) amount of fat: 0.12*H+0.0*S=0.02*(1gallon)
0.12*H=0.02
---
Solve eq 1 for S
S=1-H
-
multiply eq 2 by 100 to work with whole numbers and solve for H
12*H=2
H=1/6
--
substitute H=1/6 into S=1-H=1-1/6=5/6
--
Answer: 1/6 of a gallon of Half and half and 5/6 of a gallon of skim milk
will yield one gallon of 2% fat milk